In a​ survey,

4242​%

of the respondents stated that they talk to their pets on the telephone. A veterinarian believed this result to be too​ high, so

sheshe

randomly selected

200200

pet owners and discovered that

7878

of them spoke to their pet on the telephone. Does the veterinarian have a right to be​ skeptical? Use the

alphaαequals=0.050.05

level of significance.

Because np 0 left parenthesis 1 minus p 0 right parenthesisnp01−p0equals=nothing

greater than>

not equals≠

less than<

equals=

​10, the sample size is

less than

greater than

​5% of the population​ size, and the sample

can be reasonably assumed to be random,

cannot be reasonably assumed to be random,

is given to not be random,

is given to be random,

the requirements for testing the hypothesis

are not

are

satisfied.

​(Round to one decimal place as​ needed.)
Answer #1

np(1 - p) = 200 * 0.42 * (1 - 0.42) = 48.7

Hence,

The correct answer in blanks will be:

48.7; >; Less than; is given to be random; are